Suppose that the money supply and the nominal GDP for a hypothetical economy are $48 billion and $336 billion, respectively. If the central bank reduces the money supply by $20 billion, by how much will nominal GDP have to fall to restore equilibrium, according to the monetarist perspective?
Answer #1

Money supply = 48

Nominal GDP, Price x Real GDP = 336

This gives velocity of money V = Nominal GDP / money supply or 336 / 48 = 7

Now money supply is 48 - 20 = 28

Velocity of money should remain same at 7.

Then nominal GDP = velocity of money x money supply = 7 x 28 = 196

Hence, nominal GDP will have to fall by (336 - 196) = 140 billion to restore equilibrium, according to the

monetarist perspective.